OpenStudy (anonymous):

Match each set of ordered pairs with their rise over run. Match Term Definition negative three comma four and negative two comma one A) negative three three comma negative four and two comma one B) negative three‾fifths negative three comma four and two comma one C) negative one three comma negative four and negative two comma one D) negative five PLEASE HELP

OpenStudy (anonymous):

I

OpenStudy (anonymous):

come on now just graph it

OpenStudy (anonymous):

I give medals, PLEASE HELP

OpenStudy (anonymous):

somebody...

OpenStudy (anonymous):

@arabpride ?

OpenStudy (texaschic101):

you need to find the slopes of these points by using the slope formula : slope = (y2 - y1) / (x2 - x1) (-3,4) x1 = -3 and y1 = 4 (-2,1) x2 = -2 and y2 = 1 now we sub slope = (1 - 4) / (-2 - (-3) slope = -3/(-2 + 3) slope = -3/1 which is the same as -3 (A) (3,-4)(2,1) slope = (1 - (-4) / (2 - 3) slope = (1 + 4) / -1 slope = 5/-1 which is the same as -5 (D) (-3,4)(2,1) slope = (1 - 4) / (2 - (-3) slope = -3/(2 + 3) slope = -3/5 (B) (3,-4)(-2,1) slope = (1 - (-4) / (-2 -3) slope = (1 + 4) / -5 slope = 5/-5 which is the same as -1 (C)

OpenStudy (arabpride):

WOW @texaschic101 uve done this question before *^.^* Good Job

OpenStudy (anonymous):

@texaschic101 DANKE! :D

OpenStudy (texaschic101):

thanks....It is really not that hard....just a matter of subbing stuff in right :)

OpenStudy (anonymous):

Can I ask one more question while you're here?

OpenStudy (anonymous):

Match each equation with its equivalent equation. Match Term Definition six x equals twenty‾four A) four x equals eight x divided by six equals six B) x divided by twelve equals three x divided by two equals sixteen C) four x equals sixteen eight x equals sixteen D) x divided by eight equals four